OpenStudy (anonymous):

HELP PLEASE)..Estimate the area of the given function, f(x) = x2 + 1 using 4 inscribed rectangles on the interval [2, 4].

OpenStudy (anonymous):

@mathsmind this one first tho

OpenStudy (anonymous):

what i understand is that we have to integrate between 2 and 4

OpenStudy (anonymous):

i am having some problems it says [Math Processing Error]

OpenStudy (anonymous):

\[int_{2}^{4} x^2 dx\]

OpenStudy (anonymous):

or is it about drawing estimating the area of the parabula using blocks of squares...

OpenStudy (anonymous):

i dont know but this one is too hard :( its worth 7.0 pts im so confused, i feel like crying

OpenStudy (anonymous):

\[int_{2}^{4} x^2 dx \]

OpenStudy (anonymous):

this work is getting to hard for me

OpenStudy (anonymous):

well you have to suffere the pain of knowledge and hardship for a better future, we all went through hard times ...

OpenStudy (anonymous):

you should fight back and not let the questions beat you ...

OpenStudy (anonymous):

i think this question is saying estimate the area using rectangles, on a proper sheet

OpenStudy (anonymous):

so first of all we need to divid the interval

OpenStudy (anonymous):

what i mean by that is to divide the interval into 4 sub inscribed intervals

OpenStudy (anonymous):

let it be 2, 2.5, 3, 3.5 and 4 because they 2 and 4 are included in the interval using square brackets notation...

OpenStudy (anonymous):

okay,i'm working on it

OpenStudy (anonymous):

are you with me, this is the logic of the question from my understanding

OpenStudy (anonymous):

I am sorry that the equation editor is not functioning propely for me, probably due to browser issue, or technical problem with open study...

OpenStudy (anonymous):

A = 0.5[f(2)+f(2.5)+f(3)+f(3.5)+f(4)

OpenStudy (anonymous):

i'm looking,i'm here trying to understand

OpenStudy (anonymous):

so what is f(2)=?

OpenStudy (anonymous):

now before that the reason why i divided the intervals by a step of 0.5 is due to the fact that dx = (4-2)/4=0.5, hence 2, 2.5...etc

OpenStudy (anonymous):

let me help you on the first one f(2) = (2)^2+1=4+1=5

OpenStudy (anonymous):

what is f(3)=?

OpenStudy (anonymous):

what is f(2.5)=?

OpenStudy (anonymous):

what is f(3.5)=?

OpenStudy (anonymous):

and finally what is f(4)=?

OpenStudy (anonymous):

give me a coplue mins as i work on finish another problem

OpenStudy (anonymous):

i don't understand how you got the f(2)= part

OpenStudy (anonymous):

ok i will solve it i am replying to 287489327498732987498237 posts at the same time ... hehehe

OpenStudy (anonymous):

lol okay, i can't keep hogging you,ill be working on others as i wait.

OpenStudy (anonymous):

now young lady, let's see what your problem is...

OpenStudy (anonymous):

step one: dx = (4-2)/4 = 1/2=0.5

OpenStudy (anonymous):

step two divide the interval [2, 4] into [2, 2.5, 3, 3.5, 4]

OpenStudy (anonymous):

step three find : A = 0.5[f(2)+f(2.5)+f(3)+f(3.5)+f(4)]

OpenStudy (anonymous):

f(2) =(2)^2+1=5, f(2.5)=(2.5)^2+1= 7.25. f(3)=(3)^2+1=10, f(3.5)=(3.5)^2+1=13.25, and f(4) = (4)^2+1=17

OpenStudy (anonymous):

: A = 0.5[5+7.25+10+13.25+17]=? use the calculator

OpenStudy (anonymous):

@Ephilo are you there...?

OpenStudy (anonymous):

yes im writing it on paper

OpenStudy (anonymous):

you need to understand what we did, it is a numerical integration method ....

OpenStudy (anonymous):

i understand